OpenStudy (anonymous):

What is the least common denominator of the expression below? g^2 / 9 - g^2 + 14 + g / 24g + 8g^2 a. (3 + g) b. 7g^2 + 24g + 9 c. 8g(3 - g) d. 8g(3 + g)(3 - g) help me please :)

OpenStudy (anonymous):

well..."Common" Denominator just means that the denominators in two (or more) fractions are common, or the same.

OpenStudy (anonymous):

is it D?

OpenStudy (anonymous):

yeah i think you may be right on this one

OpenStudy (anonymous):

i got d

OpenStudy (anonymous):

alright

jimthompson5910 (jim_thompson5910):

\[\Large \frac{g^2}{9-g^2} + \frac{14+g}{24g + 8g^2}\] \[\Large \frac{g^2}{3^2-g^2} + \frac{14+g}{24g + 8g^2}\] \[\Large \frac{g^2}{(3+g)(3-g)} + \frac{14+g}{24g + 8g^2}\] \[\Large \frac{g^2}{(3+g)(3-g)} + \frac{14+g}{8g(3 + g)}\] I'll let you finish up

OpenStudy (anonymous):

8g :d?

OpenStudy (anonymous):

can i get a medal

OpenStudy (anonymous):

so A ??? im so confused now xD

jimthompson5910 (jim_thompson5910):

you were right the first time it's best to go with your gut feeling sometimes

jimthompson5910 (jim_thompson5910):

look at the factors in the denominators and focus on the unique factors

OpenStudy (anonymous):

true that

OpenStudy (anonymous):

but its least common, and 3 + g are very common lol how is it D?

jimthompson5910 (jim_thompson5910):

the unique factors are: (3+g), (3-g), and 8g multiply them to get 8g(3+g)(3-g)

jimthompson5910 (jim_thompson5910):

(3+g) is a factor, but it's not the whole LCD

OpenStudy (anonymous):

okay i see :) thanks

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

ok perve

OpenStudy (anonymous):

welcome hey im 16 what do you expect